R. A. Miech and M. J. Shanahan investigated the association of de- pression with age and education, based on a 1990 nationwide telephone survey of 2,031 adults aged 18 to 90. Of particular interest was their finding that the association of depression with education strengthens with increasing age - a phenomenon they called divergence hypoth- esis, They constructed a depression score from responses to several related questions. Education was categorized as (i) college degree, (ii) high school degree plus some college, or (11) high school degree only (a) Write down a linear regression model in which the mean depres- sion score changes linearly with age in all three categories, with possibly unequal slopes and intercepts. Identify a single parame- ter that measures the diverging gap between categories (iii) and (i) with age. If you use indicator variables, make sure you define them (b) Modify the model to specify that the slopes of the regression lines with age are equal in categories (i) and (ii) but possibly different in category (iii). Again identify a single parameter measuring divergence
Answer #1

(a)

The linear regression model is,

M δ4H * Age Bo + βι Age+ δι HC + δ2 H + δ3HC * Age+ ean Depression score

where H is categorical variable with value 1 for High School degree only (iii) and 0 for College degree (i) and High School degree plus some college (ii)

HC is categorical variable with value 1 for High School degree plus some college (ii) and 0 for College degree (i) and  High School degree only (iii)

For category (i), HC = H = 0, the linear model becomes,

Mean Depression Score-30+ β1 Age

For category (iii), HC = 0 and H = 1, the linear model becomes,

= β0+ β1 Age+ δ2+ δ4Age M ean Depression Score

-(30+δ2)+(A+δ4) Age M ean Depression core

Thus, 04 is the parameter that measure between categories (iii) and (i) with age.

(b)

Now, if the slope of the regression lines with age are equal in categories (i) and (ii), then delta_3 HC * age parameter in the previous linear model should be omitted.

Mean Depression Score-30+ β, Age+ δι HC+ δ2 H + δ4H * Age

Then for category (ii), HC = 1, H = 0

= β0+ β1 Age+ δ1 M ean Depression Score

-(30+δ1)+A Age M ean Depression core

thus, the slopes for both categories (i) and (ii) is 61

Again here,  04 is the parameter that measure between categories (iii) and (i) with age.

For category (i), HC = H = 0, the linear model becomes,

Mean Depression Score-30+ β1 Age

For category (iii), HC = 0 and H = 1, the linear model becomes,

= β0+ β1 Age+ δ2+ δ4Age M ean Depression Score

-(30+δ2)+(A+δ4) Age M ean Depression core

Thus, 04 is the parameter that measure between categories (iii) and (i) or (ii) with age.